# OffLight: An Offline Multi-Agent Reinforcement Learning Framework for Traffic Signal Control

### v1

@@ 3 OffLight Framework
OffLight is designed to tackle the challenges of offline multi-agent reinforcement learning (MARL) in the domain of traffic signal control (TSC). It combines a self-supervised learning approach with importance sampling (IS) to account for heterogeneous policies and improve learning from static, pre-collected datasets. IS and RBPS weights are collected before the offline RL training and stored. During the offline RL training phase, these weights are used to improve learning. Below, we outline the architecture and key components of OffLight, highlighting how it handles mixed-policy data to enhance traffic management performance.
@@
3.1 Dataset Structure
@@
where:
  • $\mathcal{O}_{t}=\left\{o_{t}^{1},o_{t}^{2},\dots,o_{t}^{N}\right\}$ denotes the observations at time step $t$ for each of the $N$ intersections.
  • $\mathcal{A}_{t}=\left\{a_{t}^{1},a_{t}^{2},\dots,a_{t}^{N}\right\}$ represents the joint actions taken by the agents at time $t$.
  • $\mathcal{O}^{\prime}_{t}=\left\{o_{t+1}^{1},o_{t+1}^{2},\dots,o_{t+1}^{N}\right\}$ are the subsequent observations after taking actions $\mathcal{A}_{t}$.
  • $\mathcal{R}_{t}$ is the reward received at time $t$.
  • $T_{k}$ is the length of trajectory $\tau_{k}$.
@@
This structure reflects real-world data collected from sensor-equipped intersections and enables learning traffic control policies from historical data.
3.2 Self-Supervised Learning of Behavior Policies
To effectively model diverse behavior policies in offline traffic signal control (TSC) data, OffLight uses a Gaussian Mixture Model-based Variational Graph Autoencoder (GMM-VGAE), designed to capture both spatial and temporal dependencies while addressing policy heterogeneity across intersections and time.
  • GMM for Policy Heterogeneity: The GMM structure in the VGAE models the latent space as a mixture of Gaussian distributions, allowing OffLight to represent multiple underlying policies. This disentanglement of varied traffic control behaviors within the latent space enables OffLight to more accurately capture and leverage diverse policies from real-world datasets.
  • Graph Attention Networks (GATs) and LSTM for Spatial-Temporal Dynamics: GATs capture spatial dependencies by allowing each agent to attend to relevant neighboring intersections. This attention mechanism is structured to combine local information effectively, creating a global latent representation of behavior policies across the network. Additionally, GATs enable efficient parameter sharing across agents, improving scalability for TSC in a manner similar to MARL by processing local data in a structured and computationally manageable way. LSTM layers capture the temporal evolution of traffic patterns by processing sequences of observations over time, preserving the dynamic dependencies needed for effective policy modeling.
Together, these components enable GMM-VGAE to accurately model both local and global interactions in the traffic network, ensuring OffLight’s robustness to the diverse and dynamic nature of traffic conditions. This architecture makes OffLight highly scalable and effective in learning from heterogeneous TSC datasets.
By constructing a latent space with Gaussian mixtures and capturing spatial-temporal dependencies through GAT and LSTM layers, OffLight provides a structured way to estimate behavior policies at each intersection. This enables seamless integration of importance sampling (IS) and Return-Based Prioritized Sampling (RBPS). Together, these techniques enhance learning from offline datasets by prioritizing transitions aligned with the target policy and derived from high-reward episodes. Below, we discuss how IS and RBPS are applied within OffLight to balance distributional shifts and improve sample efficiency, advancing offline reinforcement learning for traffic signal control (TSC).
Encoder: Captures spatial dependencies using Graph Attention Networks (GATs) and temporal dynamics with a Long Short Term Memory (LSTM) layer. This allows the model to effectively process both the local and global traffic behavior.
$\mathbf{z}=\text{Encoder}(\mathcal{O}_{t},\mathcal{A}_{t-1},\mathbf{A})$
Latent Space: Represents the underlying policies as a mixture of Gaussian variables, enabling the disentanglement of different policy behaviors across the network.
$p(\mathbf{z})=\sum_{k=1}^{K}\pi_{k}\mathcal{N}(\mathbf{z}\mid\mu_{k},\Sigma_{k})$
Decoder: Reconstructs the policy distribution based on the latent representation and current observation, providing an estimate of the behavior policy for each traffic signal.
$\mathcal{A}_{t}=\text{Decoder}(\mathcal{O}_{t},\mathbf{z}_{t},\mathbf{A})$
represents the estimated probability distribution $\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\cdot)$ over actions for agent $n$ at time step $t$, conditioned on the current observation and global latent policy representation.
A benefit of using GATs is that parameters can be shared across agents, thereby reducing model complexity and improving the learning efficiency. By learning these latent representations, OffLight can accurately model heterogeneous policies and ensure that the learning process accounts for diverse traffic control strategies. A detailed explanation of GMM-VGAE can be found in Appendix A.3.
3.3 Importance sampling Integration
@@
To address the challenge of distributional shift between the behavior policy and the target policy, OffLight integrates importance sampling (IS). This mechanism adjusts the influence of each transition based on its alignment with the target policy, ensuring that the learning algorithm emphasizes relevant and high-quality data.
@@
where:
  • $w_{\text{IS},t}^{k}$ is the importance sampling weight for the transition at time step $t$ in episode $k$.
@@
  • $\hat{\pi}_{b}^{i}(a_{t}^{i}\mid\mathbf{h}_{t}^{i},\mathbf{z}_{t})$ represents the estimated behavior policy probability for agent $n$, obtained from the GMM-VGAE, conditioned on it’s hidden state $\mathbf{h}_{t}^{i}$ and the latent embedding of behavior policy $\mathbf{z}_{t}$.
Importance sampling corrects for the discrepancy between the behavior policy and the target policy by re-weighting transitions  [^2] The target policy is the policy that the offline RL algorithm seeks to optimize, while the behavior policy refers to the policy used to generate the offline dataset.. Transitions where the actions are more probable under the target policy relative to the behavior policy receive higher importance sampling weights.
3.4 Return-Based Prioritized Sampling
To enhance sample efficiency and accelerate the learning process, OffLight employs Return-Based Prioritized Sampling (RBPS), similar to [27]. This strategy prioritizes episodes based on their cumulative rewards, ensuring that the learning algorithm focuses on more successful traffic control experiences.
@@ 3.4 Return-Based Prioritized Sampling
where:
  • $w_{\text{RBPS}}^{k}$ is the Return-Based Prioritized Sampling weight for episode $k$.
  • $G_{k}=\sum_{t=0}^{T_{k}}r_{t}$ is the total reward for episode $k$.
  • $G_{\text{min}}$ and $G_{\text{max}}$ are the minimum and maximum total rewards across all episodes in the dataset, respectively.
  • $p_{\text{base}}$ is a small positive constant added to ensure that all episodes have a non-zero probability of being sampled.
  • $C$ is a normalization constant to maintain numerical stability or ensure the weights sum to a desired value.
@@
3.5 Combining importance sampling and Return-Based Prioritized Sampling
@@
OffLight synergistically combines importance sampling (IS) and Return-Based Prioritized Sampling (RBPS) to address both distributional shift and sample efficiency in offline MARL for TSC. This combined weighting scheme ensures that the learning algorithm emphasizes transitions that are both aligned with the target policy and derived from high-reward episodes, thereby enhancing the effectiveness and robustness of policy learning.
@@
importance sampling adjusts the weights based on policy alignment, while RBPS prioritizes transitions from successful episodes, ensuring that the model learns from both relevant and high-quality data.  [^3] Multiplying $w_{\text{IS},t}^{k}$ and $w_{\text{RBPS}}^{k}$ ensures that only transitions that are both relevant to the target policy and come from successful episodes are given significant weight. This dual prioritization enhances the effectiveness and robustness of policy learning.
Normalization: After combining, the weights are normalized within each minibatch to maintain their probabilistic interpretation and ensure numerical stability.
$\tilde{w}_{\text{combined},t}^{k}=\frac{w_{\text{combined},t}^{k}}{\sum_{j}w_{\text{combined},t}^{j}}$ (7)
Clipping: To prevent extremely large weights from destabilizing the training process, the combined weights are clamped to a maximum value.
[…]
Integration into Loss Functions: The normalized and clamped combined weights are applied to the Temporal Difference (TD) error and the Conservative Q-Learning (CQL) loss to adjust the influence of each transition during training.
3.6 Integration with Offline RL Algorithms
OffLight can easily be integrated with existing offline RL algorithms. CQL and TD3+BC are chosen as baseline offline RL algorithms. The combined weighting scheme—comprising importance sampling (IS) weights and Return-Based Prioritized Sampling (RBPS) weights—is incorporated into both the loss functions and the sampling mechanisms of these algorithms to ensure effective bias correction and sample prioritization.
@@ 3.6 Integration with Offline RL Algorithms
By integrating these techniques, OffLight provides a robust framework for learning effective traffic control policies from static datasets, even in the presence of policy heterogeneity.
4 Experimental Setup
We conduct experiments to evaluate OffLight’s performance across three real-world urban traffic scenarios: Jinan (12 intersections), Hangzhou (16 intersections), and Manhattan (196 intersections). These scenarios vary in complexity and traffic demand, allowing for a comprehensive evaluation of OffLight’s scalability and robustness.
@@ 4.1 Traffic Scenarios
Table 1: Traffic demand levels and network size for the three urban traffic scenarios used in the experiments, including Jinan, Hangzhou, and Manhattan.
The experiments include three distinct traffic networks: Jinan, a small network with moderate traffic complexity (12 intersections); Hangzhou, a medium network with higher traffic flow (16 intersections); and Manhattan, a large urban network (196 intersections) with complex traffic patterns. Each scenario captures different levels of complexity, simulating low, medium, and high traffic demand conditions as shown in Table 1.
4.2 Data Collection
The dataset used for training includes traffic scenarios generated using a variety of controllers, including traditional rule-based methods (Fixed Time, Greedy, Max Pressure, Self-Organizing Traffic Light (SOTL)) and expert reinforcement learning controllers (MAPPO). Random policy controllers are also used to simulate suboptimal behavior. Each simulation runs for 1 hour, divided into 10 episodes of 6 minutes each, across three traffic demand levels (low, medium, and high). This results in 100 hours of simulation per controller for each demand level.
4.3 Baselines
OffLight’s performance is compared against several baseline offline reinforcement learning methods: Behavior Cloning (BC), Conservative Q-Learning (CQL), and TD3+BC. Each algorithm is trained for $20$k timesteps.
4.4 Evaluation Metrics
Performance is evaluated using two key metrics: Average Travel Time (ATT) and Queue Length (QL). These metrics are calculated across different traffic demand levels to assess the effectiveness of OffLight in managing traffic flow and reducing congestion.
5 Results and Discussion
@@
5.1 Performance on Mixed-Quality Data
We assess OffLight’s performance on datasets with varying proportions of expert and random policy data. The results are shown for three traffic networks: Jinan, Hangzhou, and Manhattan, under low, medium, and high traffic demand.
@@ 5.1 Performance on Mixed-Quality Data
Table 2: Average Travel Time (ATT): Mixed Policy Dataset
@@
Table 3: Queue Length (QL): Mixed Policy Dataset
@@
5.1.1 Average Travel Time (ATT)
Figure 5 compares the performance of OffLight with baseline offline RL methods across three traffic scenarios—Jinan, Hangzhou, and Manhattan—at low, medium, and high traffic demand levels. Lower ATT values indicate more efficient traffic flow.
Across all three scenarios (Jinan, Hangzhou, Manhattan), OffLight consistently outperforms both TD3+BC and CQL baselines. In particular, OffLight (TD3+BC) shows the best performance under high traffic demand conditions, particularly in the Jinan and Manhattan scenarios. In the Manhattan scenario, where traffic complexity is the highest, OffLight achieved up to a 6.9% reduction in ATT at high demand levels when integrated with CQL, and a 7.8% reduction when integrated with TD3+BC, as compared to their respective baselines. This highlights OffLight’s ability to handle large-scale networks effectively under congested conditions. In low traffic demand scenarios, the performance improvements were less pronounced, which could indicate that under lighter traffic, the margin for improvement is naturally smaller. The results suggest that OffLight’s incorporation of IS and RBPS allows it to prioritize transitions that are most relevant to traffic efficiency, even in datasets containing a mix of expert and suboptimal policies. By focusing learning on higher-quality data, OffLight is able to reduce congestion more effectively, particularly in high-demand, complex networks like Manhattan.
5.1.2 Queue Length (QL)
@@
Table 2 presents the queue length results for the three traffic scenarios, indicating the efficiency of OffLight in reducing the number of vehicles queued at intersections. OffLight (TD3+BC) shows significant reductions in queue lengths, particularly in medium to high traffic demand scenarios. For example, in the Hangzhou scenario with medium traffic, OffLight (TD3+BC) reduced the queue length by 11.2% compared to TD3+BC alone. In Manhattan, OffLight also showed marked reductions in queue length, outperforming both baselines across all traffic levels. For instance, OffLight (CQL) achieved a 6% reduction in queue length under high traffic demand, which is critical in reducing congestion in highly complex urban networks. However, under low traffic demand, queue length reductions are less pronounced, likely because fewer vehicles are on the road, making it easier for any control policy to manage traffic effectively. The strong performance in queue length reduction, particularly in medium and high traffic scenarios, highlights OffLight’s ability to coordinate intersections more effectively. The IS mechanism ensures that OffLight learns to prioritize actions that lead to better overall traffic management, minimizing stoppages at intersections.
Figure 6 provides a detailed visualization of the learning curves for various algorithms in the Jinan scenario on medium traffic. TD3+BC and CQL improve queue length reduction over time, with TD3+BC converging faster than CQL, but they still result in higher overall queue lengths compared to OffLight. OffLight (TD3+BC) and OffLight (CQL) both significantly outperform the other baselines, maintaining the lowest queue lengths throughout training, with OffLight (TD3+BC) performing the best overall. BC, TD3+BC, and CQL, on the other hand, show more variance and slower convergence. This indicates that OffLight is better at focusing on transitions that matter, effectively reducing congestion more consistently. Importance sampling (IS) ensures that transitions generated under different behavior policies are re-weighted based on their relevance to the current policy, correcting for any mismatch between the offline data and the target policy. This adjustment allows OffLight to more effectively utilize the dataset and focus learning on actions that lead to better traffic flow, resulting in faster convergence and significantly lower queue lengths.
5.1.3 Behavior Representation
@@
Figure 7 visualizes the t-SNE embeddings of the state-action-next state pairs and the learned latent trajectories for the Hangzhou scenario. The t-SNE visualizations reveal clear separation between the controllers in the latent space. This separation is especially evident between expert, random, and rule-based policies, with distinct clusters forming for each type of policy. This demonstrates the effectiveness of Graph-GMVAE in capturing the latent structure of the dataset.
The results also indicate that OffLight is robust to variations in the quality of the dataset. As the proportion of expert data decreases, OffLight continues to outperform baseline methods. This robustness is critical in real-world settings where data quality may vary significantly across different intersections and traffic conditions.
5.2 Ablation Studies
5.2.1 Performance Comparison with Different Levels of Mixing

### v3

@@ III OffLight Framework
OffLight is designed to tackle the challenges of offline MARL in the domain of traffic signal control (TSC). It combines a self-supervised learning approach with IS to account for heterogeneous policies and improve learning from static, pre-collected datasets. IS and RBPS weights are collected before the offline RL training and stored (see Figure 2). During the offline RL training phase, these weights are used to improve learning. Below, we outline the architecture and key components of OffLight, highlighting how it handles mixed-policy data to enhance traffic management performance.
@@
III-A OffLight Agent Design
Observation Space: Each traffic signal agent has a limited range of observation. This range reflects the realistic coverage of common sensors such as inductive loop detectors or cameras. For each incoming lane, the agent observes: (1) number of vehicles $\{n_{l}\}_{l=1}^{L_{i}}$, where $L_{i}$ is the number of lanes for traffic signal $i$, average speed of vehicles $\{s_{l}\}_{l=1}^{L_{i}}$, normalized by the speed limit, number of halted vehicles, representing the queue length $\{q_{l}\}_{l=1}^{L_{i}}$, and the current phase ID, indicating the traffic signal’s current phase. These local observations provide the sensory input required for each agent to make real-time decisions about the traffic signal phases.
Action Space: Each traffic signal agent is responsible for managing the signal phases at its respective intersection. The agent’s available actions involve selecting one of the predefined green phases applicable to that intersection. Once a green phase is chosen, the system enforces a mandatory transition to a yellow phase. The action selection occurs at fixed intervals of 5 simulation seconds.
Reward Function: We use minimizing the queue length as objective. The reward for each agent $i$ at each time step $r^{i}_{t}$ is defined as $r^{i}_{t}=-\sum_{l\in L_{i}}q_{l}(t),$ where $q_{l}(t)$ represents the queue length on lane $l$ at time $t$, and $L_{i}$ is the set of incoming lanes for traffic signal $i$. This reward structure encourages agents to minimize congestion and improve traffic flow throughout the system.
III-B Dataset Structure
@@
where $\mathcal{O}_{t}=\left\{o_{t}^{1},o_{t}^{2},\dots,o_{t}^{N}\right\}$ denotes the observations at time step $t$ for each of the $N$ intersections, $\mathcal{A}_{t}=\left\{a_{t}^{1},a_{t}^{2},\dots,a_{t}^{N}\right\}$ represents the joint actions taken by the agents at time $t$, $\mathcal{O}^{\prime}_{t}=\left\{o_{t+1}^{1},o_{t+1}^{2},\dots,o_{t+1}^{N}\right\}$ are the subsequent observations after taking actions $\mathcal{A}_{t}$, $\mathcal{R}_{t}$ is the reward received at time $t$, and $T_{k}$ is the length of trajectory $\tau_{k}$. This structure reflects real-world data collected from sensor-equipped intersections and enables learning traffic control policies from historical data.
III-C Self-Supervised Learning of Behavior Policies
OffLight employs GMM-VGAE to model heterogeneous policies, using: (1) GMM-VGAE for Policy Diversity: A Gaussian mixture graph variational autoencoder model (GMM) captures distinct traffic control policies. (2) GATs and LSTMs for Spatial-Temporal Dynamics: Graph attention networks (GATs) encode local interactions, while LSTMs track temporal trends.
These components construct a structured latent space for estimating policies at each intersection. The encoder processes agent observations and past actions to capture spatial-temporal dependencies. The latent space represents diverse traffic control policies using a GMM-VGAE model. Finally, the decoder reconstructs policy distributions based on the latent features and current observations. We also enable parameter sharing among agent, reducing model complexity while improving efficiency. This design allows OffLight to effectively model heterogeneous policies across intersections.
III-D Importance sampling Integration
@@
To address the challenge of distributional shift between the behavior policy and the target policy, OffLight integrates IS. This mechanism adjusts the influence of each transition based on its alignment with the target policy, ensuring that the learning algorithm emphasizes relevant and high-quality data.
@@
IS corrects for the discrepancy between the behavior policy and the target policy by re-weighting transitions  [^3] The target policy is the policy that the offline RL algorithm seeks to optimize, while the behavior policy refers to the policy used to generate the offline dataset.. Transitions where the actions are more probable under the target policy relative to the behavior policy receive higher importance sampling weights.
III-E Return-Based Prioritized Sampling
To enhance sample efficiency and accelerate the learning process, OffLight employs RBPS, similar to [15]. This strategy prioritizes episodes based on their cumulative rewards, ensuring that the learning algorithm focuses on more successful traffic control experiences.
@@ III-E Return-Based Prioritized Sampling
where $w_{\text{RBPS}}^{k}$ is the Return-Based Prioritized Sampling weight for episode $k$, $G_{k}=\sum_{t=0}^{T_{k}}r_{t}$ is the total reward for episode $k$, $G_{\text{min}}$ and $G_{\text{max}}$ are the minimum and maximum total rewards across all episodes in the dataset, respectively. $p_{\text{base}}$ is a small positive constant added to ensure that all episodes have a non-zero probability of being sampled, and $C$ is a normalization constant to maintain numerical stability or ensure the weights sum to a desired value.
@@
III-F Combining importance sampling and Return-Based Prioritized Sampling
@@
OffLight integrates IS and RBPS to improve sample efficiency and mitigate distributional shift in offline MARL for TSC. IS ensures transitions align with the target policy, while RBPS prioritizes high-reward episodes, refining learning by emphasizing both relevance and quality [^4] Multiplying $w_{\text{IS},t}^{k}$ and $w_{\text{RBPS}}^{k}$ ensures that only transitions that are both relevant to the target policy and come from successful episodes are given significant weight. This dual prioritization enhances the effectiveness and robustness of policy learning..
@@
To maintain stability, the combined weights are normalized and clipped within each minibatch. These adjusted weights are then incorporated into the Temporal Difference (TD) error and Conservative Q-Learning (CQL) loss functions, ensuring more effective policy learning.
III-G Integration with Offline RL Algorithms
OffLight can easily be integrated with existing offline RL algorithms. CQL and TD3+BC are chosen as baseline offline RL algorithms. The combined weighting scheme—comprising IS weights and RBPS weights—is incorporated into both the loss functions and the sampling mechanisms of these algorithms to ensure effective bias correction and sample prioritization.
@@ III-G Integration with Offline RL Algorithms
IV Experimental Setup
We evaluate OffLight in three real-world urban traffic scenarios  [^5] https://github.com/traffic-signal-control/sample-code/: Jinan (12 intersections), Hangzhou (16 intersections), and Manhattan (196 intersections). These scenarios vary in size and traffic demand, allowing for a comprehensive assessment of OffLight’s scalability and robustness. Traffic demand is classified as low, medium (real-world data), or high by adjusting the number of vehicles entering the network proportionally.
@@ IV Experimental Setup
TABLE I: Traffic demand levels and network sizes for Jinan, Hangzhou, and Manhattan. Demand represents vehicles generated per incoming lane per hour.
Traffic Scenarios: The selected networks represent different levels of complexity: Jinan (moderate, 12 intersections), Hangzhou (higher traffic, 16 intersections), and Manhattan (large-scale, 196 intersections). Each scenario simulates low, medium, and high demand conditions (Table I). Data Collection: Training data includes traffic simulations using multiple controllers: rule-based methods (Fixed Time, Greedy, Max Pressure, SOTL), expert RL controllers (MAPPO), and random policies for suboptimal behavior. Each simulation runs for 1 hour, divided into 10 episodes of 6 minutes each, across three demand levels, totaling 100 simulation hours per controller per demand level. Baselines: OffLight is compared against Behavior Cloning (BC), Conservative Q-Learning (CQL), and TD3+BC. Each algorithm is trained for 20k timesteps. Evaluation Metrics: Performance is measured using Average Travel Time (ATT) and Queue Length (QL), assessing traffic flow and congestion reduction across different demand levels.
V Results and Discussion
@@
V-A Performance on Mixed-Quality Data
We assess OffLight’s performance on datasets with varying proportions of expert and random policy data. The results are shown for three traffic networks: Jinan, Hangzhou, and Manhattan, under low, medium, and high traffic demand. Results show that OffLight consistently outperforms baselines, particularly in high-demand conditions, demonstrating its ability to manage congestion in large-scale networks.
@@ V-A Performance on Mixed-Quality Data
TABLE II: Average Travel Time (ATT): Mixed Policy Dataset
@@
TABLE III: Queue Length (QL): Mixed Policy Dataset
@@
V-A1 Average Travel Time (ATT)
Figure 5 compares OffLight to offline RL baselines. OffLight achieves lower ATT across all scenarios, with the greatest improvements in high-demand conditions. In Manhattan, OffLight reduces ATT by up to 7.8% (TD3+BC variant) and 6.9% (CQL variant) compared to their respective baselines, proving its scalability in complex networks. Performance gains are smaller in low-traffic scenarios, where traffic flow is already manageable. These results indicate that OffLight effectively prioritizes high-quality data, improving traffic efficiency in congested environments.
V-A2 Queue Length (QL)
@@
Table III and Figure 6 show that OffLight significantly reduces queue lengths, especially in medium and high traffic conditions. In Hangzhou, OffLight (TD3+BC) decreases queue lengths by 11.2% compared to TD3+BC alone. In Manhattan, OffLight (CQL) achieves a 6% reduction under high traffic demand, demonstrating its capability to optimize traffic flow at intersections. Improvements are smaller in low-traffic scenarios where congestion is naturally less severe.
Figure 6 highlights the learning curves for the Jinan scenario under medium traffic demand. OffLight (TD3+BC) and OffLight (CQL) achieve faster convergence and lower queue lengths compared to baselines, with OffLight (TD3+BC) performing best. TD3+BC and CQL improve over time but maintain higher queue lengths than OffLight, while BC and other baselines show greater variance and slower convergence.
OffLight’s IS mechanism ensures that transitions are weighted based on their relevance to the learned policy, reducing reliance on low-quality data. This re-weighting helps OffLight make better traffic control decisions, resulting in smoother traffic flow, quicker convergence, and significantly lower queue lengths across all demand levels.
V-A3 Behavior Representation
@@
Figure 7 visualizes t-SNE embeddings of learned latent space in the Hangzhou scenario, illustrating OffLight’s ability to differentiate between expert, random, and rule-based policies. The distinct clusters confirm that GMM-VGAE effectively captures structural differences in control policies, enabling more robust policy learning.
OffLight remains effective even as the proportion of expert data decreases, consistently outperforming baselines. This robustness is crucial in real-world settings where dataset quality varies across intersections and traffic conditions, ensuring reliable performance despite heterogeneous data sources.
V-B Ablation Studies
V-B1 Performance Comparison with Different Levels of Mixing